OpenStudy (anonymous):

Damasko asked Sophia to choose a number between 1 and 20. He told her to add 7 to that number, multiply by 4, subtract 6 and divide by 2. Sophia’s final number was 35. What was her original number?

OpenStudy (mathstudent55):

Call the unknown number x. Add 7 to the number: x + 7 Multiply by 4: 4(x + 7) Subtract 6: 4(x + 7) - 6 Divide by 2: [4(x + 7) - 6]/2 Result equals 35: [4(x + 7) - 6]/2 = 35 Solve the equation below for x: \(\dfrac{4(x + 7) - 6}{2} = 35 \)

OpenStudy (anonymous):

Okay, I'll try it.

OpenStudy (mathstudent55):

Start by multiplying both sides by 35. That will get rid of the denominator on the left side.

OpenStudy (mathstudent55):

Any luck?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

I got 12

OpenStudy (mathstudent55):

That is it. Good job! To check, just do all the operations listed in the problem and make sure you get 35 as the result.

OpenStudy (anonymous):

Okay, thanks!

OpenStudy (mathstudent55):

wlcm